x-y-z=0
x-2y-2z=3
-2x+2y-z=3

Answer by checkley77(12569)

x-y-z=0 multiply this equation by 2 & add to the third equation.
x-2y-2z=3
**********************
-2x+2y-z=3
2x-2y-2z=0 now add.
----------------------
-3z=3
z=3/-3
z=-1 answer.
x-y+1=0
x-2y+6=3 now multiply this equation by -1 & add to the previous equation.
**********************************
x-y+1=0
-x+2y-6=-3 now add.
--------------------
y-5=-3
y=-3+5
y=2 answer.
x-2+1=0
x-1=0
x=1 answer.